OpenStudy (anonymous):

Which equation would best help solve this problem? Five added to 3 times a number is equal to 8 less than five times the number.

OpenStudy (mathstudent55):

Are you given choices?

OpenStudy (anonymous):

Yes @mathstudent55

OpenStudy (anonymous):

3x + 5 = 5x – 8 3x – 4 = 5x + 8 4x + 3 = 8x – 5 8x + 5 = 4x + 3

OpenStudy (anonymous):

3 x + 5 = 5 x – 8 3 x – 4 = 5 x + 8 4 x + 3 = 8 x – 5 8 x + 5 = 4 x + 3

OpenStudy (anonymous):

@mathstudent55 idk how to get rid of those diamonds.

OpenStudy (mathstudent55):

I see the equations.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

They are using x as the variable to represent the unknown number.

OpenStudy (mathstudent55):

If x is the unknown number, how do you represent 3 times the number?

OpenStudy (anonymous):

You multiply it I think We had a session on this but i want there.

OpenStudy (mathstudent55):

Right. x is the number. 3 times the number is 3 times x which is written as 3x Ok?

OpenStudy (anonymous):

Ohhh ok.

OpenStudy (mathstudent55):

Now we need to add 5 to that product. "Five added to 3 times a number" is: 3x + 5 Ok so far?

OpenStudy (anonymous):

I'm Getting it so far.

OpenStudy (mathstudent55):

The next word is "is". In math, "is" usually means = We now have 3x + 5 =

OpenStudy (mathstudent55):

In fact, you have "is equal to" which for sure is "="

OpenStudy (mathstudent55):

Now let's look at what comes after "is equal to"

OpenStudy (anonymous):

And then it says 8 less times

OpenStudy (mathstudent55):

8 less than five times the number.

OpenStudy (mathstudent55):

Let's start with 5 times the number. Just like 3 times the number was 3x, what is 5 timers the number?

OpenStudy (anonymous):

5x

OpenStudy (mathstudent55):

Good. "less" means you are taking away, so it means subtraction. 8 less than 5x means take away 8 from 5x That means subtract 8 from 5x How do you think you'd write that using a subtraction sign?

OpenStudy (anonymous):

8-5x

OpenStudy (anonymous):

or was it 5x-8

OpenStudy (mathstudent55):

No. You need to be careful. The way you wrote it, you are taking away 5x from 8. We need to take 8 away from 5x.

OpenStudy (mathstudent55):

Exactly. Now you got it.

OpenStudy (anonymous):

YA!

OpenStudy (mathstudent55):

Now we can fill in our equation: 3x + 5 = 5x - 8

OpenStudy (mathstudent55):

One of the choices is exactly that equation.

OpenStudy (anonymous):

Thank You so MUCH! I would have never figured out this question withour u!

OpenStudy (mathstudent55):

You are welcome. Here is a helpful hint. When you see a problem like this, break up the sentence into small parts. Work on each part. Then put it all together.

OpenStudy (mathstudent55):

Remember the keywords: lees means subtraction added is obviously addition is means =

OpenStudy (anonymous):

And i have one more if its not to much and its kinda like this last one but a little more confusing.

OpenStudy (mathstudent55):

Yes, but please start a new post.

OpenStudy (mathstudent55):

I'll look for it.

OpenStudy (anonymous):

ok!